OpenStudy (amistre64):

find: f''(x) , when f(x) = \(2e^x-3sin(x)+2x\cfrac{1-e^x)}{pi}-2\)

OpenStudy (amistre64):

ack.. i left a " ) " hanging in there

OpenStudy (angela210793):

how come ur asking the others? O.o

OpenStudy (amistre64):

becasue it breaks up the algebra blahdom

OpenStudy (amistre64):

i should have asked to go from the f'' to f with conditions of f(0) = f(pi) = 0

OpenStudy (amistre64):

\[f''(x)=2e^x +3sin(x)\] was the original ; given that f(0) = f(pi) = 0

OpenStudy (amistre64):

and that is what i came up with :)

OpenStudy (angela210793):

ahhhhaaaaa..well let me think

OpenStudy (amistre64):

gotta get for a few, class starting. Have fun with it and ill check on yall later

OpenStudy (angela210793):

kk,see ya :D

OpenStudy (angela210793):

My opinion(not sure it is right though) u have to integrate to find f'(x) which is 2e^x-3cosx+c then integrate 2e^x-3cosx+c and u'll get cx+2e^x-3sinx+c then find c since u have f(0)

OpenStudy (anonymous):

I am Getting \(f"(x) = 2e^x - \frac{e^x}{\pi} +3sinx\)

OpenStudy (zarkon):

\[f(x)=2e^x-3\sin(x)+2x\cfrac{1-e^\pi}{\pi}-2\]

OpenStudy (zarkon):

\[e^\pi\text{ not }e^x\] for the second 'e' in the function

OpenStudy (anonymous):

ah Lol sorry I am wrong

OpenStudy (angela210793):

how did u guys get tht?

OpenStudy (anonymous):

yea you got me... Thanks @Zarkon

OpenStudy (anonymous):

hi \[\cos 30\]

OpenStudy (amistre64):

yeah, i got that right on the paper but wrong on the latex :) e^pi thnx zarkon